OpenStudy (anonymous):

Suppose an object moves in an elliptical orbit parametrized by x = 2 cos t and y = 7 sin t, where t ≥ 0. Determine the value of the velocity at the indicated times: v(pi/4) v(pi/2)

OpenStudy (anonymous):

for v(pi/4) I got 2 (which was wrong) by taking the sqrt((-2sin(pi/4))^2) + (2cos(pi/4))^2)).. ifyou have any idea where I went wrong, I would appreciate the help!!

OpenStudy (anonymous):

Velocity is a vector.

OpenStudy (anonymous):

I don't think you need to find the magnitude.

OpenStudy (anonymous):

\[ v_x=-2\sin t\\ v_y=7\cos t \]

OpenStudy (anonymous):

now to find magn. take sqrt of ((-2sint)^2+(7cost)^2)

OpenStudy (anonymous):

subst. t =pi/4 or pi/2